# Little By Little: Continual Learning via Self-Activated Sparse Mixture-of-Rank Adaptive Learning

### v2

@@ 4 Experiments
In this paper, we conduct experiments across a diverse set of tasks, including continual learning for both vision-language CLIP models and LMs, and analyze catastrophic forgetting during fine‐tuning. Detailed implementation settings and more experiment results are provided in Appendix A.1.
CL of CLIP. We evaluate on two benchmarks using CLIP-ViT/B-16 (Radford et al., 2021; Ilharco et al., 2021) on MTIL (Zheng et al., 2023; Yu et al., 2024) and X-TAIL (Xu et al., 2024; Lu et al., 2024), with a range of image classification datasets. Following prior work, we report Transfer (zero-shot on unseen tasks), Last (retention on earlier tasks), and Average (mean accuracy across all datasets of all learning tasks) to assess continual learning performance.
CL of LM. We follow prior works (Qin & Joty, 2021; Razdaibiedina et al., 2023; Wang et al., 2023; Qiao & Mahdavi, 2024) to continually fine-tune T5-large (Raffel et al., 2020) and LLaMA2-7B (Touvron et al., 2023) on five text-classification benchmarks, under three different task orderings. We report the average final accuracy across all tasks after completing the last task.
Generalization and forgetting on unseen tasks. To assess effects on pre-trained general knowledge, we fine-tune Llama3.1-8B (Grattafiori et al., 2024) on the CodeAlpaca code-generation dataset (Chaudhary, 2023) and evaluate zero-shot in-domain performance on HumanEval (Chen et al., 2021), as well as out-of-domain accuracy on a broad selection of MMLU (Hendrycks et al., 2021) subjects.
Table 2: Zero-shot in-domain performance on HumanEval (Pass@1) for the code generation domain and out-of-domain accuracy on selected MMLU subjects (formal logic, philosophy, world religions, economics, public relations, STEM, physics, machine learning) after fine-tuning Llama-3.1-8B on CodeAlpaca. The last two columns report trainable parameters (for MoRA: added / activated).
Method
HumanEval
(Pass@1)
Out-of-Domain (Acc.)
Params
@@
MoRA 47.56 48.41 70.09 82.46 73.53 68.18 55.53 46.08 41.96 63.70 41.9/26.2 0.52%/0.32%
4.1 Continual Learning of CLIP
Table 1 reports X-TAIL results (MTIL results in Appendix Table 11). Prior methods (Yu et al., 2024; Xu et al., 2024) preserve pre-trained capabilities via zero-shot predictions and are therefore capped by the base model’s zero-shot performance. Following recent practice (Zheng et al., 2023; Lu et al., 2024), MoRA continuously trains the pre-trained model so it can surpass that zero-shot ceiling and better generalize to unseen domains, while still preserving existing capabilities. Across both Last and Average metrics, MoRA substantially outperforms prior works. Unlike methods that rely on domain-ID prediction (Yu et al., 2024), high-dimensional projection layers or feature-banks (Xu et al., 2024), MoRA improves the pre-trained representation without external mechanisms; compared with (Lu et al., 2024), which learns fixed weights per rank-$r$ update, MoRA uses input-dependent gates on individual rank-one updates for finer expert specialization and dynamic utilization.
Table 3: Summary of results on standard CL benchmarks with T5-large. We report averaged accuracy after training on the last task across three task orderings.
Standard CL Benchmark
Method Order-1 Order-2 Order-3 Avg.
MTL 80.0
SeqFT 18.9 24.9 41.7 28.5
SeqLoRA 44.6 32.7 53.7 43.7
IncLoRA 66.0 64.9 68.3 66.4
Replay 55.2 56.9 61.3 57.8
EWC 48.7 47.7 54.5 50.3
LwF 54.4 53.1 49.6 52.3
L2P 60.3 61.7 61.1 60.7
LFPT5 67.6 72.6 77.9 72.7
InfLoRA 75.2 75.4 75.8 75.5
O-LoRA 75.4 75.7 76.3 75.8
LB-CL 76.9 76.5 76.8 76.7
MoRA 77.4 77.5 77.9 77.6
4.2 Continual Learning of LM
Table 3 reports results across three task orderings: MoRA consistently outperforms prior methods and closely approaches the multi-task learning (MTL) upper bound. Unlike O-LoRA (Wang et al., 2023) and LB-CL (Qiao & Mahdavi, 2024), which rely on orthogonality constraints or gradient projections between per-task LoRA adapters (potentially limiting adapter capacity), MoRA needs no extra regularization. By decomposing each rank-r update into rank-one components and applying self-activated, sparse gating, MoRA lets each component specialize on its own input distribution, reducing interference and more effectively capturing diverse patterns. We further evaluate MoRA on LLaMA2-7B (Touvron et al., 2023) under the same continual-learning setup (Table 9 in the Appendix), MoRA also outperforms O-LoRA by 2.3% averaged over 3 task orders.
4.3 Forgetting and Generalization during Fine‐Tuning
Table 2 examines how standard fine-tuning affects both in-domain performance and out-of-domain generalization. We observe two key phenomena. (1) Forgetting of unrelated knowledge: conventional LoRA fine-tuning causes drops on topics that are unrelated to the fine-tuning data (e.g., Philosophy, World Religions, Public Relations), indicating that pre-trained general knowledge is overwritten during adaptation. (2) Improved generalization: fine-tuning on code data increases accuracy on logically and quantitatively oriented tasks (Formal Logic, Economics, STEM, Physics, Machine Learning), showing that representations learned from code transfer to related symbolic domains.
In contrast, MoRA both preserves unrelated pre-trained knowledge and improves in- and out-of-domain generalization. Decomposing each low-rank update into specialized rank-one components and activating only the most relevant subset prevents catastrophic overwriting of prior knowledge while allowing targeted adaptation. At the same time, MoRA achieves strong code-generation performance using roughly one-third of the activated parameters required by a rank-32 LoRA and substantially improves out-of-domain accuracy.
4.4 Visualizations of Fine-grained Rank Activations
Figure 3 shows rank activations recorded during continual learning. These visualizations illustrate two properties of MoRA: (1) individual ranks specialize on distinct input patterns, and (2) the model substantially reduces cross-task interference, thereby mitigating forgetting. Extended visualizations across more tasks and scenarios appear in Fig. 4 and Fig. 5 in the Appendix.
Each rank specializes in distinct input patterns. In Fig. 3(a), airplane patches strongly activate Rank 0, while blue-sky backgrounds predominantly activate Rank 11. In Fig. 3(c), Ranks 19, 20, and 29 (orange boxes) respond to jaguar patches. The more complex backgrounds in Fig. 3(c) (e.g., leaves, shadows) yield a richer, more distributed pattern than the simple blue sky in Fig. 3(a), highlighting MoRA’s capacity to model contextual complexity. Some ranks learned earlier are also reused on later tasks (Fig. 5, Appendix), indicating transfer of shared semantics.
Reduced cross-task interference and mitigated forgetting. Comparing the same input after Task 1 (Fig. 3(a)) and after Task 2 (Fig. 3(b)) shows almost identical activations (more in Fig. 4, Appendix): Rank 0 still responds to airplane semantics and Rank 11 to blue-sky patches. This stability indicates that our self-activated, sparse mixture-of-ranks prevents later updates from overwriting earlier task representations, reducing inter-task interference and mitigating catastrophic forgetting.
@@ 4.4 Visualizations of Fine-grained Rank Activations
(a) Rank Activations of MoRA on data from Task 1 after learning Task 1.
@@
(b) Rank Activations of MoRA on data from Task 1 after learning Task 2.
@@
(c) Rank Activations of MoRA on data from Task 2 after learning Task 2.
Figure 3: Visualization of MoRA rank activations during Task 1 and Task 2 training. Activations are extracted from the K projection in the attention module (layer 8) of the image encoder. Corresponding image patches are shown below each activation map, with regions relevant to each class marked by orange bounding boxes. Zoom in for details. More visualizations are in Figs. 4 and 5 of the Appendix, demonstrating forgetting mitigation and knowledge reuse.
4.5 Ablation Studies
Routing/activation strategy. Table 4 compares different strategies for computing mixture weights, under a controlled setup with one LoRA per task, standard training loss, and no additional regularization: (1) MoE-LoRA trains a router to assign weights to each LoRA adapter; its coarse-grained design and lack of mechanisms that encourage expert specialization lead to greater interference and increased forgetting compared to MoRA. (2) Non-self-activation trains a finer-grained router for mixing each rank-one component. However, with a much larger number of experts (ranks), the router becomes increasingly prone to confusion, resulting in higher training complexity and degraded performance compared with MoE-LoRA. (3) Self-activation infers weights directly from softmax-normalized activations (Eq. 8), matching or exceeding Non-self-activation and demonstrating that each rank can infer its own relevance without extra trainable router modules. (4) Top-k rank selection restricts updates to the k most relevant ranks, reducing interference and improving accuracy. (5) Sharpness control concentrates activation mass and substantially boosts specialization. (6) Test-time rank thresholding removes low-importance ranks at inference, yielding a slight gain by eliminating noisy contributions.
Table 4: Routing/activation strategy
Mixture Calculation Transfer Average Last
MoE-LoRA 62.56 69.45 74.53
Non-self-activation 60.09 65.97 69.76
Self-activation 60.26 65.94 69.85
    w/ top-k rank selection 60.69 66.52 70.62
    w/ sharpness control with $\tau_{\text{MoRA}}$ 62.07 71.15 79.62
    w/ test-time rank thresholding 60.78 66.83 71.08
    MoRA 63.30 72.70 80.90
@@ 4.5 Ablation Studies
(b) Temp. $\tau_{\text{MoRA}}$
@@
Figure 3: Ablation on (a) rank activation budget, (b) temperature $\tau_{\text{MoRA}}$, and (c) threshold $\delta$.
Rank activation budget. We probe how many ranks are necessary by varying the activation budget (Fig. 4(a)). To isolate this effect, we disable other sparsity mechanisms (e.g., thresholding). After training on 10 tasks (160 ranks total), increasing the budget from 2 (1.25%) to 16 active ranks (10%) yields a large improvement in final accuracy and then gradually plateaus. This suggests that each rank specializes in distinct input patterns and that a small, sparse subset of ranks is sufficient to cover diverse inputs. Even with a high budget, the self-activating gating remains robust: it still privileges the most relevant ranks per token while suppressing lower-relevance contributions.
Sharpness control of $\tau_{\text{MoRA}}$. We use temperature scaling (Eq. (10)) to control mixture sharpness: lower $\tau_{\text{MoRA}}$ yields sharper activations, while higher $\tau_{\text{MoRA}}$ flattens the distribution and engages more ranks. Fig. 4(b) shows that increasing $\tau_{\text{MoRA}}$ improves transfer to unseen tasks, even surpassing the reported state-of-the-art Transfer in Table 1 when $\tau_{\text{MoRA}}$ is high. A moderate $\tau_{\text{MoRA}}$ around 0.01 balances specificity and generalization, delivering strong Last and Average scores.
@@ 5 Conclusion
We present MoRA, a self-activated sparse mixture-of-ranks framework that decomposes each LoRA update into rank‐1 experts and self‐infer their sparse mixture weights. By reducing knowledge redundancy and resolving router–expert mismatch during CL, MoRA delivers superior continual‐learning performance, significantly reduces forgetting, and enhances generalization.
Future works. MoRA controls mixture sharpness and sparsity via top-k selection, temperature scaling, and thresholding. While effective, the sparse activation mechanism can be further improved by incorporating awareness of the input data distribution. Learning this temperature or adapting the mixture’s sharpness based on data offers a promising avenue for future research. We also aim to extend the method’s applicability to a broader range of PTMs and applications.
Reproducibility Statement
Further implementation details and dataset descriptions are provided in Sec. 4 and Appendix A.1. Our source code will be publicly released upon paper acceptance.

### v3

@@ 4 Experiments
In this paper, we conduct experiments across a diverse set of tasks, including continual learning for both vision-language CLIP models and LLMs, and analyze catastrophic forgetting during fine‐tuning. Detailed implementation settings and more experiment results are provided in Appendix A.1.
4.1 Experimental Results
Continual Learning of CLIP. We evaluate X-TAIL performance in Table 1, reporting Transfer, Last, and Average accuracy (MTIL results in Appendix Table 12). While prior methods [96, 91] are constrained by the base model’s zero-shot ceiling, MoRAM follows [105, 50] to continuously adapt, surpassing this limit and achieving superior Last and Average scores. Critically, MoRAM enhances representation without the external domain-IDs or feature banks required by [96, 91], and improves upon the fixed weighting in [50] by utilizing dynamic, input-dependent gating for finer expert specialization.
Continual Learning of LLMs. In Table 2, we evaluate MoRAM on LLaMA [23] and Gemma [71] using the TRACE benchmark [81]. Traditional regularization [36], rehearsal [48], and prompt-based [85] methods struggle to scale, often underperforming simple In-Context Learning. Among LoRA variants, naive SeqLoRA suffers significant forgetting, while recent approaches like O-LoRA [80], HiDeLoRA [79], and TreeLoRA [60] rely on rigid orthogonality or complex structural expansions. In contrast, MoRAM performs robustly on LLMs, achieving results competitive with existing methods. We provide additional evaluations on language classification tasks in Tables 8 and 9 of the Appendix.
Table 2: Comparison with a broad range of CL methods on the TRACE benchmark. We report Overall Performance (OP (%) $\uparrow$) and Backward Transfer (BWT (%) $\downarrow$). Results are averaged over three runs with standard deviations. The best results are highlighted in bold.
FIX(ICL) SeqLoRA OGD GEM EWC L2P DualPrompt HiDeLoRA O-LoRA TreeLoRA MoRAM
meta-llama / LLaMA-2-7B-Chat
OP 38.94 $\pm$ 0.3 34.3 $\pm$ 1.2 42.09 $\pm$ 1.6 40.08 $\pm$ 1.6 42.36 $\pm$ 1.2 36.23 $\pm$ 0.8 37.69 $\pm$ 1.2 41.60 $\pm$ 0.8 42.78 $\pm$ 0.8 43.52 $\pm$ 1.0 44.54 $\pm$ 0.9
BWT 18.5 $\pm$ 0.8 8.06 $\pm$ 1.2 6.77 $\pm$ 1.2 5.97 $\pm$ 0.8 8.25 $\pm$ 0.8 8.03 $\pm$ 0.8 7.12 $\pm$ 0.4 7.16 $\pm$ 0.4 3.46 $\pm$ 0.4 1.37 $\pm$ 0.3
google / Gemma-2B-it
OP 32.3 $\pm$ 0.2 31.89 $\pm$ 0.8 32.85 $\pm$ 1.4 26.48 $\pm$ 1.5 28.35 $\pm$ 1.6 31.14 $\pm$ 1.2 32.42 $\pm$ 1.0 33.25 $\pm$ 0.9 33.73 $\pm$ 0.8 33.41 $\pm$ 0.9 36.27 $\pm$ 0.7
BWT 15.28 $\pm$ 0.4 12.27 $\pm$ 0.9 18.25 $\pm$ 0.9 16.96 $\pm$ 1.2 15.77 $\pm$ 0.7 14.25 $\pm$ 0.5 13.66 $\pm$ 0.5 12.36 $\pm$ 0.4 8.50 $\pm$ 0.5 2.74 $\pm$ 0.4
meta-llama / LLaMA-3-1B-Instruct
OP 31.16 $\pm$ 0.4 29.73 $\pm$ 1.6 30.12 $\pm$ 2.0 32.19 $\pm$ 2.0 31.96 $\pm$ 1.6 29.38 $\pm$ 1.2 30.76 $\pm$ 1.2 33.73 $\pm$ 1.2 32.94 $\pm$ 0.8 36.14 $\pm$ 0.7 37.77 $\pm$ 0.8
BWT 17.03 $\pm$ 1.2 15.2 $\pm$ 1.6 10.74 $\pm$ 1.6 11.62 $\pm$ 1.2 13.57 $\pm$ 0.8 11.34 $\pm$ 0.8 12.36 $\pm$ 0.8 12.89 $\pm$ 1.2 7.36 $\pm$ 0.8 3.12 $\pm$ 0.8
Table 3: Standard fine-tuning of Llama-3.1-8B on CodeAlpaca. We report zero-shot in-domain performance on HumanEval (Pass@1) for the code generation and out-of-domain accuracy on selected MMLU subjects (formal logic, philosophy, world religions, economics, public relations, STEM, physics, machine learning). The last two columns report trainable parameters (for MoRAM: added / activated).
@@ 4.1 Experimental Results
MoRAM 47.56 48.41 70.09 82.46 73.53 68.18 55.53 46.08 41.96 63.70 41.9/26.2 0.52%/0.32%
Forgetting and generalization after standard fine‐tuning. Table 3 examines how standard fine-tuning affects both in-domain performance and out-of-domain generalization. (1) Forgetting: Dense LoRA updates overwrite pre-trained knowledge, degrading performance on semantically distant topics (e.g., Religions). (2) Positive transfer: Conversely, code adaptation improves logically related tasks (e.g., STEM, Logic), indicating valid transfer potential.
In contrast, MoRAM leverages its atomic memory structure, treating updates as independent rank-1 memory atoms rather than entangled blocks, the model isolates new knowledge accumulation from existing representations. This granular independence prevents the catastrophic overwriting of unrelated concepts (e.g., Religions) while the content-addressable self-activation ensures that relevant atoms are reused for logical tasks (e.g., STEM, Logic). Consequently, MoRAM achieves superior out-of-domain accuracy using roughly one-third of the active parameters required by a standard rank-32 LoRA.
@@
(a) Rank Activations of MoRAM on data from Task 1 after learning Task 1.
@@
(b) Rank Activations of MoRAM on data from Task 1 after learning Task 2.
@@
(c) Rank Activations of MoRAM on data from Task 2 after learning Task 2.
Figure 3: Visualization of MoRAM rank activations during Task 1 and Task 2 training. Activations are extracted from the K projection in the attention module (layer 8) of the image encoder. Corresponding image patches are shown below each activation map, with regions relevant to each class marked by orange bounding boxes. Zoom in for details. More visualizations are in Figs. 5 and 6 of the Appendix, demonstrating forgetting mitigation and knowledge reuse.
Table 4: Routing Strategies.
Routing Transfer Average Last
MoE-LoRA (Coarse Baseline) 62.56 69.45 74.53
External Router (Learned $W_{r}$) 60.09 65.97 69.76
Self-Activated Retrieval 60.26 65.94 69.85
    w/ Sparsity Constraint (Top-$k$) 60.69 66.52 70.62
    w/ Temperature Scaling ($\tau_{\text{MoRAM}}$) 62.07 71.15 79.62
    w/ Threshold-based expert selection ($\delta$) 60.78 66.83 71.08
    MoRAM (Full) 63.30 72.70 80.90
@@
(b) Temp. $\tau_{\text{MoRAM}}$
@@
Figure 3: Ablation on (a) rank activation budget, (b) temperature $\tau_{\text{MoRAM}}$, and (c) threshold $\delta$.
4.2 Visualizations of Rank-1 Memory Activations
Figure 3 shows rank activations recorded during continual learning. These visualizations illustrate two properties of MoRAM: (1) individual ranks specialize on distinct input patterns, and (2) the model substantially reduces cross-task interference, thereby mitigating forgetting. Extended visualizations across more tasks and scenarios appear in Fig. 5 and Fig. 6 in the Appendix.
Each rank specializes in distinct input patterns. In Fig. 3(a), airplane patches strongly activate Rank 0, while blue-sky backgrounds predominantly activate Rank 11. In Fig. 3(c), Ranks 19, 20, and 29 (orange boxes) respond to jaguar patches. The more complex backgrounds in Fig. 3(c) (e.g., leaves, shadows) yield a richer, more distributed pattern than the simple blue sky in Fig. 3(a), highlighting MoRAM ’s capacity to model contextual complexity. Some ranks learned earlier are also reused on later tasks (Fig. 6, Appendix), indicating transfer of shared semantics.
Reduced interference and mitigated forgetting. Comparing the same input after Task 1 (Fig. 3(a)) and after Task 2 (Fig. 3(b)) shows almost identical activations (more in Fig. 5, Appendix): Rank 0 still responds to airplane semantics and Rank 11 to blue-sky patches. This stability indicates that our self-activated, sparse mixture-of-ranks prevents later updates from overwriting earlier task representations, reducing interference and mitigating forgetting.
4.3 Ablation Studies
Ablation of memory retrieval strategies. Table 4 analyzes the impact of different retrieval mechanisms under a controlled setup. (1) MoE-LoRA (Coarse Baseline) employs a router to weight entire adapters. Its coarse granularity forces the simultaneous activation of conflicting subspaces, leading to significant interference and forgetting. (2) External Router applies a separately learned router to individual rank-1 atoms. However, decoupling the routing logic from the memory content leads to retrieval collapse: as the number of atoms grows, the external router struggles to index them precisely, resulting in worse performance than the coarse baseline. (3) Self-Activated Retrieval replaces the external router with intrinsic key-value matching (Eq. 5). By ensuring the retrieval condition is strictly aligned with the expert’s content, it resolves the routing ambiguity without extra parameters. (4) Sparsity Constraint (Top-$k$) enforces a strict activation budget. This serves as a gating mechanism that prevents memory collisions and reduces interference during training. (5) Temperature Scaling sharpens the retrieval distribution. This concentrates gradient flow on the most relevant “specialist” atoms, accelerating their adaptation while protecting shared memories. (6) Threshold-based selection prunes weak activation signals ($\delta$) strictly at test time, maximizing retrieval precision by eliminating noise.
Rank Activation Budget. We probe rank usage by varying the activation budget on a 10-task checkpoint (Fig. 4(a)). To isolate this effect, we disable other sparsity mechanisms (e.g., thresholding). Performance improves as the budget increases from 2 to 16 (10% of total ranks) before plateauing, confirming that a small, sparse subset of specialized ranks suffices to cover diverse inputs. Crucially, accuracy remains robust even at high budgets, demonstrating that self-activation naturally prioritizes relevant signals while suppressing noise, even without strict sparsity constraints.
Sharpness enhancement. Temperature scaling (Eq. (7)) modulates activation sparsity: lower $\tau_{\text{MoRAM}}$ concentrates probability mass, while higher $\tau_{\text{MoRAM}}$ broadens expert participation. Fig. 4(b) demonstrates that higher $\tau_{\text{MoRAM}}$ significantly boosts transfer to unseen tasks (even surpassing SOTA in Table 1). However, we select a moderate $\tau_{\text{MoRAM}}$ of 0.01 to optimally balance specificity (retention) and generalization (transfer), ensuring strong overall performance.
@@ 5 Conclusion
We presented MoRAM, a framework that challenges the coarse-grained design of standard MoE-LoRA. By redefining adaptation as the accumulation of atomic associative memories, we demonstrated that high-performing CL does not require complex external routers, but rather precise, content-addressable retrieval of low-rank atomic memories.
Future works. MoRAM controls mixture sharpness and sparsity via top-k selection, temperature scaling, and thresholding. While effective, the sparse activation mechanism can be further improved by incorporating awareness of the input data distribution. Learning this temperature or adapting the mixture’s sharpness based on data offers a promising avenue for future research. We also aim to extend the method’s applicability to a broader range of PTMs and applications.
Impact Statement
MoRAM’s sparse mixture-of-ranks method makes continual adaptation of large vision–language and language models far more efficient and reliable. By updating only the most relevant low-rank subspaces, it reduces compute and memory requirements, enabling on-device personalization in domains like education, healthcare, and finance. This democratizes access to powerful, continually evolved models for smaller teams and resource-constrained settings, driving innovation and improving quality of life. Preserving pre-trained knowledge also ensures stability in safety-critical systems, such as medical diagnostics and autonomous vehicles, where forgetting of pre-trained capabilities can be dangerous.
The focus of this paper is fundamental research, and is broadly applicable to model fine-tuning techniques. Thus, it is possible that MoRAM’s lightweight updates could be misused to insert stealthy backdoors or reinforce hidden biases that are hard to detect. Because it leaves most pre-trained parameters untouched, existing biases may persist or become entrenched. To mitigate these risks, we recommend strict audit logging of rank-level updates, anomaly detection on sparse changes, robust access controls, and routine bias and fairness assessments alongside any deployment.
Acknowledgments and Disclosure of Funding
This work was partially supported by the ARC DECRA Fellowship (DE230101591) awarded to D. Gong. H. Lu is affiliated with CSIRO Data61 through a PhD scholarship and acknowledges the support of the Google PhD Fellowship.
@@ Appendix A Additional Details and Results of Experiments
A.1 Detailed Experiment Settings
Continual Learning of CLIP on X-TAIL and MTIL. The MTIL setting consists of 1,201 classes drawn from 11 diverse datasets: Aircraft [53], Caltech101 [25], CIFAR100 [39], DTD [14], EuroSAT [31], Flowers [58], Food [7], MNIST [18], OxfordPet [59], Cars [38], and SUN397 [90]. In the X-TAIL setting, a total of 10 datasets are used, with CIFAR100 [25] excluded to prevent domain overlap, following the protocol in [91]. In line with [91], we use a 5-shot split for MTIL and a 16-shot split for X-TAIL.
We follow the experimental setups in [105, 96, 91] and use the CLIP model with a ViT-B/16 backbone [63] for all experiments. By default, MoRAM is applied to every pre-trained weight matrix in both the vision and text encoders, with an initial rank of 16 per update. Each task is trained for 500 iterations using AdamW [49] with a learning rate of 5e45e-4. During continual learning, we freeze the ranks learned from previous tasks and initialize new $r=16$ ranks for each incoming task. The rank activation budget is set to 16 throughout all tasks. We set the temperature $\tau_{\text{MoRAM}}=0.1$ and the threshold $\delta=0.2$.
Continual Learning of LLMs on TRACE. We evaluate MoRAM on the TRACE benchmark [81], a comprehensive suite designed to assess continual learning in LLMs such as LLaMA [72] and Gemma [71]. TRACE standardizes eight diverse datasets spanning domain-specific understanding (C-STANCE, FOMC), reasoning (ScienceQA, NumGLUE-cm/ds), summarization (MeetingBank), code generation (Py150), and multilingual simplification (20Minuten). Following standard protocols, each task consists of 5,000 training and 2,000 testing examples. We employ task-specific metrics to capture performance nuances: Accuracy for classification and reasoning tasks; ROUGE-L for summarization; SARI for simplification; and similarity scores for code generation.
We adopt the variable epoch schedule from Qian et al. [60] ({2, 1, 3, 2, 1, 2, 2, 3} epochs). For each incoming task, we initialize $r=16$ new ranks while maintaining a fixed rank activation budget of $k=16$ across all tasks. We set the softmax temperature $\tau_{\text{MoRAM}}=0.03$ and the inference threshold $\delta=0.2$. Optimization proceeds with a learning rate of $1\times 10^{-4}$ and a batch size of 4, utilizing a maximum context length of 1024 tokens. All experiments leverage DeepSpeed ZeRO-2 with BF16 mixed-precision on a cluster of four Nvidia H100 GPUs.
Continual Learning of LMs on language classification tasks. We follow the protocol of previous work in continually fine-tuning the T5-large [64] and LLaMA2-7B [72] on a suite of text-classification tasks. We train on five standard benchmarks—AG News, Amazon Reviews, Yelp Reviews, DBpedia, and Yahoo Answers—using three distinct task orderings drawn from [62, 65, 80, 61]. To probe longer sequences, we extend this to a 15-dataset stream (Table 5), incorporating tasks from the original CL benchmark [102], GLUE [75], SuperGLUE [74], and the IMDB movie reviews corpus. Natural language prompts for each task are presented in Table 6, with NLI tasks (MNLI, RTE, CB), sentiment classification (Amazon, Yelp, SST-2, IMDB), and topic classification (AG News, DBpedia, Yahoo).
We evaluate three distinct task sequences for both the standard CL and 15-task benchmarks (Table 7). After completing the final task in each stream, we report the average accuracy across all tasks. All experiments use one epoch per task with DeepSpeed, a fixed learning rate of 1e31e-3, batch size 64, and dropout of 0.1. MoRAM is applied to both the query and key projection matrices within attention layers, initializing $r=8$ new ranks for each incoming task similarly as in [80, 61]. We maintain a constant activation budget of 4 ranks throughout continual learning, set the temperature $\tau_{\text{MoRAM}}=0.1$, and the threshold $\delta=0.2$.
Generalization and forgetting on unseen tasks after standard fine-tuning. To assess effects on pre-trained general knowledge, we fine-tune Llama3.1-8B [28] on the CodeAlpaca code-generation dataset [8] using llama-Factory [104] and evaluate using lm-eval-harness [26] on zero-shot in-domain performance on HumanEval [11], as well as out-of-domain accuracy on a broad selection of MMLU [32] subjects—Formal Logic, Philosophy, World Religions, Economics, Public Relations, STEM, Physics, and Machine Learning.
In this experiment, MoRAM is applied to all linear weight matrices of the pre-trained model. We fine-tune on CodeAlpaca with a batch size of 32 over 3 epochs and a cosine learning-rate schedule starting at 5e45e-4. We train with $r=16$ ranks and enforce a constant activation budget of 4 ranks. The self-routed gating uses a temperature $\tau_{\text{MoRAM}}=0.5$ and a threshold $\delta=0.2$. We observe that, due to variations in hidden representations across architectures, the optimal temperature setting can differ across each pre-trained models.